OpenStudy (anonymous):

find the solution to the initial value problem. dy/dx=x((x^2)+1)/9y^3 y(0)=-1/sqrt(3)

OpenStudy (zzr0ck3r):

can you get all the x's on one side and all the y's on one side?

OpenStudy (zzr0ck3r):

\[\frac{dy}{dx}=\frac{x(x^{2}+1)}{9y^{3}}\rightarrow 9y^{3}*dy=x(x^{2}+1)dx\]

OpenStudy (zzr0ck3r):

you with me?

OpenStudy (anonymous):

yea

OpenStudy (zzr0ck3r):

\[\int9y^{3}dy=\int x(x^{2}+1)dx\]

OpenStudy (zzr0ck3r):

so \[\int 9y^{3}dy=\int (x^{3}+x)dx\]

OpenStudy (zzr0ck3r):

can you solve this?

OpenStudy (anonymous):

9/4y^(4)=(x^(4)/4)+(x^(2)/2)+c

OpenStudy (zzr0ck3r):

right now solve for y

OpenStudy (anonymous):

y=((x^(4)/9)+(2x^(2)/9)+(4/9c))^(1/4)

OpenStudy (zzr0ck3r):

now use the initial condition to solve for c

OpenStudy (zzr0ck3r):

so plug in 0 for x, and -1/sqrt(3) for y

OpenStudy (anonymous):

is c=-1/4?

OpenStudy (zzr0ck3r):

use 9/4y^(4)=(x^(4)/4)+(x^(2)/2)+c plug in the point (0,-1/sqrt(3)) 9/4(-1/sqrt(3))^(4)=(0^(4)/4)+(0^(2)/2)+c c=1/4

OpenStudy (zzr0ck3r):

the negative goes away

OpenStudy (anonymous):

oh right. careless mistake

OpenStudy (anonymous):

is it ok if i send you two other problems as an attached file and can you help me look over them because i dont know where my error is?

OpenStudy (zzr0ck3r):

sure, do you know if we got this one right?

OpenStudy (zzr0ck3r):

I trusted your integration so it should be fine..

OpenStudy (anonymous):

im going to type in the answer right now give me a moment :)

OpenStudy (anonymous):

i got it wrong. i'll attach my work and show you.

OpenStudy (zzr0ck3r):

the one we just did?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

hmm let me write it on paper real fast, I hate using the latek here I am no good...

OpenStudy (anonymous):

i'll be back in 5 minutes

OpenStudy (zzr0ck3r):

\[y=\sqrt[4]{\frac{x^{4}+2x^{2}+1}{9}}\] is this what you got?

OpenStudy (anonymous):

i had a 4/9c at the end is that wrong? (i attached the file of my work)

OpenStudy (zzr0ck3r):

sorry on phone, give me a few minutes...my wife is lost lol

OpenStudy (anonymous):

yeah its fine

OpenStudy (anonymous):

wait nevermind. yes i got that answer you got at the end

OpenStudy (zzr0ck3r):

oh ok, good. :)

OpenStudy (zzr0ck3r):

need anything else?

OpenStudy (anonymous):

yes just one problem. i'll attach it.

OpenStudy (anonymous):

actually im still working on it right now. i actually got it though thank you for your help :)

OpenStudy (zzr0ck3r):

good deal One thing to look out for. say we have 3y=3x+c and we divide by 3 y=x+c/3 many people think (because we can do this with normal integration for a general solution), that we can just say y=x+c because after all c is a constant, don't do this. Keep c separated and then solve for c when you have a ivp. You didn't have this problem but it is something I just remembered.